# Evolution and learning in differentiable robots

### v1

@@ Abstract
The automatic design of robots has existed for 30 years but has been constricted by serial non-differentiable design evaluations, premature convergence to simple bodies or clumsy behaviors, and a lack of sim2real transfer to physical machines. Thus, here we employ massively-parallel differentiable simulations to rapidly and simultaneously optimize individual neural control of behavior across a large population of candidate body plans and return a fitness score for each design based on the performance of its fully optimized behavior. Non-differentiable changes to the mechanical structure of each robot in the population—mutations that rearrange, combine, add, or remove body parts—were applied by a genetic algorithm in an outer loop of search, generating a continuous flow of novel morphologies with highly-coordinated and graceful behaviors honed by gradient descent. This enabled the exploration of several orders-of-magnitude more designs than all previous methods, despite the fact that robots here have the potential to be much more complex, in terms of number of independent motors, than those in prior studies. We found that evolution reliably produces “increasingly differentiable” robots: body plans that smooth the loss landscape in which learning operates and thereby provide better training paths toward performant behaviors. Finally, one of the highly differentiable morphologies discovered in simulation was realized as a physical robot and shown to retain its optimized behavior. This provides a cyberphysical platform to investigate the relationship between evolution and learning in biological systems and broadens our understanding of how a robot’s physical structure can influence the ability to train policies for it. Videos and code at sites.google.com/u.northwestern.edu/eldir.
@@
Figure 2: A diversity of body shapes were discovered by evolution, each with its own unique internal distribution of active (green) and passive (gray) springs. The optimized performance (displacement in the desired direction of travel) achieved by gradient based learning is denoted above each design in meters (m). On rugged terrains bipedal body plans often emerged (A2, A6, B1, B4, B5, D3, E4, F3, F5) whereas on flat terrain robots often evolved three or four legs (A3, C1, C3, C4). Occasionally, swinging limbs that did not make contact with the ground were used to generate forward momentum (A4, B1, B4, E5). Videos and code can be found at sites.google.com/u.northwestern.edu/eldir.
@@ II Methods
Figure 8: Sim2real. A design discovered in simulation (A-D) locomotes over half a body length forward after 5 actuation cycles (D). A robot manufactured with the same design (E-H) achieved a similar displacement after 5 actuation cycles (H). To determine whether the robot’s movement was a result of chance or due to the robot’s evolved design, the design was altered by replacing one of its two motors with a passive spring (I-L). The physical instantiation of the evolved design moves in the desired direction of travel (green arrow in D and H), but, when one of its motors was replaced with a passive spring, it moved the “wrong way” (i.e. in reverse direction; red arrow in L). The physical trajectories (light purple traces) and the simulated trajectories (dark purple traces) of a set of masses near the actuating springs—those along the posterior edge of the design (m1-m5)—were tracked during behavior (M). Vertical range of motion (N) at these parts of the physical robot’s body was relatively consistent with the simulator’s predictions: the most movement occurred closest to the actuating springs (m3 and m4) and the least occurred at the mass in contact with the surface plane (O). Videos at sites.google.com/u.northwestern.edu/eldir.
@@
We occasionally observed numerical instability during learning or in the resulting mass-spring system. When sorting by fitness value, robots with loss values from any training iteration equaling to NaN (not a number) or infinity were considered invalid and discarded. To account for unstable (chaotic) mass-spring system dynamics we also filtered out robots exhibiting large absolute difference in loss value from one iteration to the next. Empirically, we found large deviations to be indicative of a control policy resulting in unstable body mechanics, which could generate large, artificial body displacement scores. Across all conducted evolutionary runs this filtering criteria eliminated less than 0.21% of the robots evaluated and the average proportion of discarded robots in each run was 0.20% ($\pm$ 0.003% SD).
@@
Parts were printed on a Nexa3D XiP resin printer with x45 Clear resin (see Table IV in Appx. Sect. VII for resin properties). After printing, parts were washed for two minutes while attached to the build plate in xCLEAN and subsequently one minute in IPA after being detached from the build plate. Parts were dried and then post cured for 25 minutes.
@@ IV Discussion
  • Problem: There is an extensive set (below we describe two examples) of limitations related to accurate simulations of physical reality and effective transfer of simulated designs and behavior to physical robots. The simulator was selected primarily for convenience of modeling the system of interest with automatic differentiation support, so the body plans and control policies discovered in simulation may not be well-suited let alone precisely tuned to exhibit the desired behavior as physical robots.
  • Solution: An iterative process of physical system characterization and simulation development could result in a tighter coupling and more effective sim2real transfer, potentially altering the search landscapes in which learning and evolution operate, which could in turn yield a completely different set of optimized designs.
  • Problem: Shape memory alloy springs—those used for prototyping—were not able to actuate with the same strength as simulated springs. In addition, simulated springs could expand or contract at any rate, however the actuation speed of the physical springs were limited in reality by physical constraints such as heating and cooling capacities.
  • Solution: Simulations could be modified to more closely match the capabilities of the physical system by disallowing instantaneous expansion and modulating the maximum actuation strength.
  • Problem: There were no limitations on inertial effects in simulation.
  • Solution: Inertia was less relevant in the behavior of larger (more massive) bodies with fewer actuators, such as the transferred design in Fig. 8.

### v2

@@ Abstract
The automatic design of robots has existed for 30 years but has been constricted by serial non-differentiable design evaluations, premature convergence to simple bodies or clumsy behaviors, and a lack of sim2real transfer to physical machines. Thus, here we employ massively-parallel differentiable simulations to rapidly and simultaneously optimize individual neural control of behavior across a large population of candidate body plans and return a fitness score for each design based on the performance of its fully optimized behavior. Non-differentiable changes to the mechanical structure of each robot in the population—mutations that rearrange, combine, add, or remove body parts—were applied by a genetic algorithm in an outer loop of search, generating a continuous flow of novel morphologies with highly-coordinated and graceful behaviors honed by gradient descent. This enabled the exploration of several orders-of-magnitude more designs than all previous methods, despite the fact that robots here have the potential to be much more complex, in terms of number of independent motors, than those in prior studies. We found that evolution reliably produces “increasingly differentiable” robots: body plans that smooth the loss landscape in which learning operates and thereby provide better training paths toward performant behaviors. Finally, one of the highly differentiable morphologies discovered in simulation was realized as a physical robot and shown to retain its optimized behavior. This provides a cyberphysical platform to investigate the relationship between evolution and learning in biological systems and broadens our understanding of how a robot’s physical structure can influence the ability to train policies for it. Videos and code at https://sites.google.com/view/eldir.
@@
Figure 2: A diversity of body shapes were discovered by evolution, each with its own unique internal distribution of active (green) and passive (gray) springs. The optimized performance (displacement in the desired direction of travel) achieved by gradient based learning is denoted above each design in meters (m). On rugged terrains bipedal body plans often emerged (A2, A6, B1, B4, B5, D3, E4, F3, F5) whereas on flat terrain robots often evolved three or four legs (A3, C1, C3, C4). Occasionally, swinging limbs that did not make contact with the ground were used to generate forward momentum (A4, B1, B4, E5). Videos and code can be found at https://sites.google.com/view/eldir.
@@ II Methods
Figure 8: Sim2real. A design discovered in simulation (A-D) locomotes over half a body length forward after 5 actuation cycles (D). A robot manufactured with the same design (E-H) achieved a similar displacement after 5 actuation cycles (H). To determine whether the robot’s movement was a result of chance or due to the robot’s evolved design, the design was altered by replacing one of its two motors with a passive spring (I-L). The physical instantiation of the evolved design moves in the desired direction of travel (green arrow in D and H), but, when one of its motors was replaced with a passive spring, it moved the “wrong way” (i.e. in reverse direction; red arrow in L). The physical trajectories (light purple traces) and the simulated trajectories (dark purple traces) of a set of masses near the actuating springs—those along the posterior edge of the design (m1-m5)—were tracked during behavior (M). Vertical range of motion (N) at these parts of the physical robot’s body was relatively consistent with the simulator’s predictions: the most movement occurred closest to the actuating springs (m3 and m4) and the least occurred at the mass in contact with the surface plane (O). Videos at https://sites.google.com/view/eldir.
@@
We occasionally observed numerical instability during learning or in the resulting mass-spring system. When sorting by fitness value, robots with loss values from any training iteration equaling to NaN (not a number) or infinity were considered invalid and discarded. To account for unstable (chaotic) mass-spring system dynamics we also filtered out robots exhibiting large absolute difference in loss value from one iteration to the next. Empirically, we found large deviations to be indicative of a control policy resulting in unstable body mechanics, which could generate large, artificial body displacement scores. Across all conducted evolutionary runs this filtering criteria eliminated less than 0.83% of the robots evaluated and the average percentage of discarded robots in each run was 0.20% ($\pm$ 0.29% SD).
@@
Parts were printed on a Nexa3D XiP resin printer with x45 Clear resin (see Table IV for resin properties). After printing, parts were washed for two minutes while attached to the build plate in xCLEAN and subsequently one minute in IPA after being detached from the build plate. Parts were dried and then post cured for 25 minutes.
@@ III Results
Figure 11: Mean fitness across the population. The average initial fitness (light green) and average optimized fitness (dark green) across all 10K designs in the current population is shown at every generation ($\pm$1 S.D.) for the five mutation-only evolutionary trials conducted on flat terrain. Unlike Figs. 5, 6, and 10, where the fitness of the most performant design is shown at every generation, here it is demonstrated the population at large contains increasingly good learners and is increasingly fit overall.
@@
TABLE IV: Material properties of the Nexa3D x45 Clear resin used to print the structure of the physical robots.
ASTM code Property Value
D638 Tensile Modulus 1600 MPa
D638 Ultimate Tensile Strength 52 MPa
D638 Tensile Elongation at Break 12%
D790 Flex Modulus 2100 MPa
D790 Flex Strength 95 MPa
D256 Notched Izod 19 J/m
D570 Water Absorption 6%
D2240 Hardness Shore D 85
@@ IV Discussion
  • Problem: Physical robots were not able to actuate with the same strength and speed as their simulated counterparts.
  • Solution: An iterative process of physical system characterization and simulation development could result in a tighter cyber-physical coupling.